If f(x)=2x^2-7x+3, find f(2).
SHOW WORK!

Answer by checkley77(12569)

F(2)=2*2^2-7*2+3
F(2)=2*4-14+3
F(2)=8-14+3
F(2)=-3 ANSWER.